coherence sml=341.8 5th harmonic of 1709
DAS profile VSOP_HO
L-band Rx is cooling
trombone is 53 steps out from IN limit
RX attenuator is 7 dB
need to check Rx attenuator once receiver is fully cool.
Rx chan 2 FET3 LNA current was zero yesterday. fixed by thermal cycling to clr frost